What is the molality of a solution containing 11.7 g Na 2 SO 4 dissolved in 2383.6 g of water?


1
Expert's answer
2021-04-26T09:06:18-0400

Сm = n / m

n = m / M

M (Na2SO4) = 142.04 g/mol

n (Na2SO4) = 11.7 / 142.04 = 0.08 mol

Cm (Na2SO4) = 0.08 / 2.3836 = 0.03 m